A farmer has a fence surrounding a 36 foot square plot and wants to use the fence to enclose a circular flower bed. The area enclosed would be?
Answer by oberobic(2304)

The square plot is 36 sq. ft., so the sides are 6 ft.
Therefore, the fence is 4*6 = 24 ft long.
In turn, the circumference of the circular plot would be 24 ft.
pi*d = 24
d = 24/pi = 7.63943727
r = 7.63943727/2 = 3.81971863
Area = pi*r^2 = 45.8366236 sq. ft.